OpenStudy (rizags):

If f is differentiable at 1, and the limit of f(1+h)/h as h->0 is 5, what is f(1)?

OpenStudy (rizags):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

` f is differentiable at 1` so \[\Large \lim_{h\to 0}\frac{f(1+h)-f(1)}{h}\] exists and it is defined based on the limit definition of the derivative

OpenStudy (rizags):

so \[f'(1) = 5-\lim_{h \rightarrow 0}\frac{f(1)}{h}\]

jimthompson5910 (jim_thompson5910):

\[\Large \lim_{h\to 0}\frac{f(1+h)}{h}=5\] \[\Large \lim_{h\to 0}\frac{f(1+h){\LARGE \color{red}{-0}}}{h}=5\] \[\Large \lim_{h\to 0}\frac{f(1+h){\LARGE \color{red}{-f(1)}}}{h}=5\] so we see that \(\Large f(1) = 0\)

OpenStudy (rizags):

so my step would be invalid?

jimthompson5910 (jim_thompson5910):

it's valid, I just don't see where to go with it

OpenStudy (rizags):

oh ok. And you knew to put in that f(1) as 0 simply because the equation resembled the derivative limit?

jimthompson5910 (jim_thompson5910):

I saw it match up with the limit definition. Well almost match up. It was just missing the `-f(1)` part

OpenStudy (rizags):

so f(x) = 0 and f'(x) = 5?

jimthompson5910 (jim_thompson5910):

f ' (1) = 5 and f(1) = 0 we can't say anything about f(x) or f ' (x) in general

OpenStudy (rizags):

ok, thank you

jimthompson5910 (jim_thompson5910):

IF it was \[\Large \lim_{h\to 0}\frac{f(x+h)-f(x)}{h}=5\] then we can say f ' (x) = 5

OpenStudy (rizags):

ok got it